OpenStudy (anonymous):

Write the equation of the circle whose center is at the origin and whose diameter is 12.

OpenStudy (anonymous):

x^2 + y^2 = r^2 is the general formula for a circle with the center in [0,0] Thus, \[x^2 + y^2 = 12^2\] \[x^2 + y^2 = 144\]

OpenStudy (anonymous):

\[x^2+y^2+z^2=36\]

OpenStudy (anonymous):

OR \[x^2+y^2=36\]

OpenStudy (unklerhaukus):

the general form a a circle centered at \((a,b)\)\[(x-a)^2+(y-b)^2=r^2\]

OpenStudy (phi):

This was (a little bit) of a trick question. They tell you the diameter is 12. But you need the radius, which is 1/2 of the diameter. The radius is 6